the measure of the supplement of an angle is 40 less than 3 times its complement, find the measure of the angle, its supplement and its complement, please give me the formula and how you figured it out. thanks so much.
Answer by nerdybill(7384)

the measure of the supplement of an angle is 40 less than 3 times its complement, find the measure of the angle, its supplement and its complement, please give me the formula and how you figured it out. thanks so much.
.
Let x = the angle
then
90-x = the complement
180-x = the supplement
.
From: "the supplement of an angle is 40 less than 3 times its complement" we get our equation:
180-x = 3(90-x) - 40
180-x = 270-3x - 40
180-x = 230-3x
180+2x = 230
2x = 50
x = 25 degrees (the angle)
.
The complement:
90-x = 90-25 = 65 degrees
.
The supplement:
180-x = 180-25 = 155 degrees